OpenStudy (anonymous):

6 * 4^x = 99 Solve for x, round to two decimal places.

OpenStudy (anonymous):

\[4^x=\frac{99}{6}\]So,\[\log_{4} 4^x=\log_{4}\frac{33}{2}\]Or,\[x=\log_{4} \frac{33}{2}=2.02\]